# DeepCRE: Revolutionizing Drug R&D with Cutting-Edge Computational Models

### 06 Mar 2024 v1

@@ Results
Some intriguing insights emerge from the drug testing results, shedding light on the deeper mechanism exploration of drug candidates with similar MoAs. For instance, the DRC of GSK1059615 exhibits a significant increase at the highest concentration compared to other PI3K/MTOR inhibitors (Fig. 6d). Perturbation data from CMAP L1000[34] database provides some explanations (Fig. 6e-f). Cells treated with GSK1059615 show up-regulated cellular survival and stress response related pathways (isomerase activity, intramolecular transferase activity and phosphotransferases) and genes (IGFBP3, MYCBP, GTF2A2), DNA repair and cytoskeleton regulation related genes (PARP1, RUVBL1, STAMBP), and down-regulated cell cycle checkpoints related pathways (protein tyrosine/MAP kinase/MAP kinase tyrosine phosphatase activity) and cell cycle related genes (STX1A, PARP6, KDM3A). The up-regulation and down-regulation of these genes and pathways may reflect the defensive mechanisms adopted by cells to resist the toxicity of high concentrations of drugs, potentially leading to the pausing of the cell cycle or the inhibition of cell death pathways[35, 36]. The other example is that the DRC of the drug Sorafenib, rather than another RTK inhibitor Dasatinib, declines at higher concentrations in three CRC organoids (Fig. 6g). It has been reported that some transport proteins may play a role in the accumulation and resistance of drugs, with Dasatinib being primarily restricted by ABCG1, and Sorafenib being mainly limited by ABCG2[37]. Further investigation suggests that ABCG2 expression, significantly lower in CRC tumor tissues compared to normal tissues[38] (Fig. 6h), may contribute to the reduced resistance of Sorafenib in CRC. Molecular docking analysis corroborates these findings, demonstrating a favorable binding affinity of Sorafenib to the target ABCG2 (Fig. 6i). Similar conclusions of binding poses and interactions have been drawn by a published paper, which presents single-particle cryo-EM studies of ABCG2 binding with another RTK inhibitor imatinib[39]. Furthermore, survival analysis reveales a negative correlation between ABCG2 expression and disease-free survival (DFS), implicating ABCG2-mediated multi-drug resistance (MDR) in CRC (Fig. 6j).
@@
a, DRC across four CRC organoids, colored by the drug set, indicating that the drug set A possesses significantly greater effectiveness than drug set C. b, Heatmap of correlation coefficients of drugs, demonstrating that the drug efficacy is related to the drug MoA and the drug set. c, Boxplot of correlation coefficients of drugs, indicating differences in the distribution of efficient populations among drugs targeting DNA replication. d, DRC of PI3K/MTOR inhibitor drugs across four CRC organoids, showing the dramatic increase in DRC of the drug GSK1059615 at the highest concentration. e, Umap plot of gene expression of cells after different perturbations, with representative differential genes shown in the dashed frame. f, Bar plot of enriched pathways, with up-regulated pathways colored in red and down-regulated pathways in blue. g, DRC of RTK inhibitor drugs across three CRC organoids, showing the decline in DRC of the drug Dasatinib, rather than Sorafenib, at higher concentrations. h, Boxplot of genes ABCG1 and ABCG2 in the TCGA-CRC cohort, indicating significantly lower expression of ABCG2 in tumor samples. i, 3D and 2D diagram of ABCG2-Sorafenib docking. j, Survival curve of ABCG2 in the disease-free survival (DFS) of TCGA-CRC corhort.

### 09 Mar 2024 v2

@@ Results
Some intriguing insights emerge from the drug testing results, shedding light on the deeper mechanism exploration of drug candidates with similar MoAs. For instance, the DRC of GSK1059615 exhibits a significant increase at the highest concentration compared to other PI3K/mTOR inhibitors (Fig. 6d). Perturbation data from CMAP L1000[35] database provides some explanations (Fig. 6e-f). Cells treated with GSK1059615 show up-regulated cellular survival and stress response related pathways (isomerase activity, intramolecular transferase activity, and phosphotransferases) and genes (IGFBP3, MYCBP, GTF2A2), DNA repair, and cytoskeleton regulation related genes (PARP1, RUVBL1, STAMBP), and down-regulated cell cycle checkpoints related pathways (protein tyrosine/MAP kinase/MAP kinase tyrosine phosphatase activity) and cell cycle related genes (STX1A, PARP6, KDM3A). The up-regulation and down-regulation of these genes and pathways may reflect the defensive mechanisms adopted by cells to resist the toxicity of high concentrations of drugs, potentially leading to the pausing of the cell cycle or the inhibition of cell death pathways[36, 37]. The other example is that the DRC of the drug Sorafenib, rather than another RTK inhibitor Dasatinib, declines at higher concentrations in three CRC organoids (Fig. 6g). It has been reported that some transport proteins may play a role in the accumulation and resistance of drugs, with Dasatinib being primarily restricted by ABCG1, while Sorafenib being mainly limited by ABCG2[38]. Further investigation suggests that ABCG2 expression, significantly lower in CRC tumor tissues compared to normal tissues[39] (Fig. 6h), may contribute to the reduced resistance of Sorafenib in CRC. Molecular docking analysis corroborates these findings, demonstrating a favorable binding affinity of Sorafenib to the target ABCG2 (Fig. 6i). Similar conclusions of binding poses and interactions have been drawn by a published paper, which presents single-particle cryo-EM studies of ABCG2 binding with another RTK inhibitor imatinib[40]. Furthermore, survival analysis reveals a negative correlation between ABCG2 expression and disease-free survival (DFS), implicating ABCG2-mediated multi-drug resistance (MDR) in CRC (Fig. 6j).
@@
a, DRC across four CRC organoids, colored by the drug set, indicating that drug set A possesses significantly greater effectiveness than drug set C. b, Heatmap of correlation coefficients of drugs, demonstrating that the drug efficacy is related to the drug MoA and the drug set. c, Boxplot of correlation coefficients of drugs, indicating differences in the distribution of efficient populations among drugs targeting DNA replication. d, DRC of PI3K/mTOR inhibitor drugs across four CRC organoids, showing the dramatic increase in DRC of the drug GSK1059615 at the highest concentration. e, Umap plot of gene expression of cells after different perturbations, with representative differential genes shown in the dashed frame. f, Bar plot of enriched pathways, with up-regulated pathways colored in red and down-regulated pathways in blue. g, DRC of RTK inhibitor drugs across three CRC organoids, showing the decline in DRC of the drug Dasatinib, rather than Sorafenib, at higher concentrations. h, Boxplot of genes ABCG1 and ABCG2 in the TCGA-CRC cohort, indicating significantly lower expression of ABCG2 in tumor samples. i, 3D and 2D diagram of ABCG2-Sorafenib docking. j, Survival curve of ABCG2 in the disease-free survival (DFS) of the TCGA-CRC cohort.
*** DISCUSSION ***
Discussion
In this study, we have proposed the DeepCRE model and demonstrated its potential to revolutionize the drug R&D process. Through advancements in pretraining strategies, DeepCRE surpasses the existing SOTA models, achieving an average performance improvement of 17.7% in patient-level CRE, and a remarkable 5-fold increase in indication-level CRE. Moreover, DeepCRE has identified six drug candidates that show significantly greater effectiveness than a comparator set of two approved drugs in 5/8 CRC organoids tested. Besides, we have uncovered some pharmacodynamic and pharmacological insights into certain drug candidates, such as GSK1059615 and Dasatinib, which could inform further R&D process. Importantly, it is not merely about identifying one or two potential drugs, rather, it highlights the capability of DeepCRE to uncover a collection of drug candidates with enhanced therapeutic effects.
There are several future directions to explore. Firstly, leveraging Large Language Models (LLMs) for patient GEPs holds tremendous potential[41, 42]. Encoding GEPs using specialized genomic or multi-modal LLMs, which integrate gene language with clinical indicators and patient descriptions, could yield more comprehensive insights. Secondly, we advocate for expanding drug R&D beyond the traditional screening paradigm to include the generation paradigm. Generating models enables the active design of molecules with desired properties, offering benefits such as targeted design, cost and time savings, and fostering innovation by creating novel chemical entities. This approach reduces reliance on extensive screening of existing compounds and holds promise for developing breakthrough therapies with higher efficacy and lower failure rates during clinical trials. Finally, establishing a synergistic dry-wet-loop framework that integrates computational predictions with experimental validations can accelerate the drug R&D process. This integration facilitates seamless translation from pre-clinical research to clinical trials, ultimately benefiting patients in need.
Methods
Construction of the patient-level CRE dataset
The patient-level CRE dataset involved the aggregation of both unlabeled and labeled data from public sources. The unlabeled dataset comprises annotations, such as tumor types and gene expression profiles sourced from the CCLE[43] and UCSC Xena[44] databases. All gene expression data was standardized using the transcripts per million (TPM) metric and subsequently log-transformed to ensure uniformity in data scale and distribution. The labeled dataset integrates drug response metrics from the GDSC database[45] and patients from the TCGA database[39]. In the case of cell lines, drug response is quantified via the area under the drug response curve (AUC), offering a comprehensive measure of drug efficacy. For patient data, the progression-free survival (PFS) period metric is utilized, reflecting the duration of response before clinical relapse[39].
Data curation involved several quality control measures. Only samples with both tumor type annotations and corresponding gene expression data were considered for inclusion. Labeled data underwent further refinement, requiring the following criteria: (i) the drug was a small molecule with a readily available SMILES representation in PubChem[46] or Drugbank[27], and (ii) the treatment regimen consists of the single-drug therapy throughout the entire duration. Tumor types with sufficient amounts of unlabeled data (more than 50 samples for patients and more than 30 samples for cell lines) and labeled data (at least seven samples for patients with more than one drug overlapping in patients and cell lines) were retained for analysis (Supplementary Fig. 3). Finally, the drug response labels were binarized based on the median response within each tumor type to facilitate analysis.
Establishment of the DeepCRE model zoo
The DeepCRE model zoo comprises eight models, including the initial autoencoder (AE), domain separation network (DSN), and their variants. Additional previous SOTA models, such as P-SDL and C-MDL models, are described in the Supplementary materials, along with detailed experiment setup information.
The initial autoencoder-based DeepCRE model
As shown in Supplementary Fig. 1d, the DeepCRE models take drug and patient representations as input. Drug embeddings are generated using a pretrained Graph Neural Network[47] (GNN) with a hidden representation dimension of 300. Patient embeddings are generated by an autoencoder designed for the initial DeepCRE model. The autoencoder is pretrained using unlabeled data from both cell lines and patients. It is then fine-tuned using labeled data from cell lines and used for zero-shot testing on patients. The encoders and decoder in the autoencoder are three-layer neural network modules with dimensions (N, 512), (512, 256), (256, 128) and (128, 256), (256, 512), (512, N), respectively, and employ the rectified linear activation function. N is the size of GEPs input, determined by tumor types and pretraining strategy. Details are described in the Supplementary materials. The classifier network used for fine-tuning consists of three-layer neural networks with dimensions (428, 64), (64, 32) using rectified linear activation, and (32, 1) using sigmoid activation.
The DSN model and its variants
The DSN model and its variants also utilize an autoencoder backbone for unlabeled training. The key difference is that the DSN model incorporates two encoders with orthogonal embeddings. The underlying assumption is that GEPs consist of both drug-response-related signals and other confounding factors, which can be encoded by the shared and private encoders, respectively. Moreover, the distributional difference of the latent embeddings between the two domains (cell lines and patients) is minimized through maximum mean discrepancy (MMD) and adversarial loss. These losses can be applied either to the concatenation of private and shared embeddings or only to the shared embeddings, corresponding to the domain separation and reconstruction network (DSRN) and DSN models, respectively. In total, the DeepCRE model zoo encompasses eight models with various training losses. For mathematical details, please refer to the Supplementary materials.
The tumor-type adaptive pretraining strategy
The tumor-type adaptive pretraining strategy is proposed to enhance the performance of DeepCRE models. Considering the significant pan-cancer heterogeneity in patient GEPs, not all patient data is suitable for pretraining a specific tumor type. tumor-type adaptive pretraining is considered more effective than the original all-data pretraining, particularly for adv-loss-based DeepCRE models (Fig. 2d). For instance, before zero-shot testing on drug response data for BRCA patients, the GEPs of both BRCA patients and all cell lines are utilized for unlabeled pretraining, followed by fine-tuning using cell line drug response data with the overlap drugs. Finally, this model is employed for zero-shot testing on the BRCA patient drug response dataset.
To demonstrate the impact of pretraining strategies on DeepCRE models, t-SNE plots are employed to visualize the distribution of embeddings for both cell lines and patients. The pretrained encoders from different strategies (test-pairwise or all-data) and methods (AE, DSN, DSN-mmd, DSN-adv) are saved and used to generate latent embeddings. These embeddings are then used for the t-SNE plot, where colors distinguish the sample types (Fig. 3d). The degree to which two colors mix together represents the alignment effect between the two domains. Two metrics, namely MMD and KL divergence, are employed to assess the alignment effect quantitatively.
Validation of DeepCRE for pharmaceutical value assessment
To validate the potential of DeepCRE in clinical pharmaceutical value assessment, the DSN-adv DeepCRE model, trained with all-drug data, was employed to screen 233 small molecules for patients across 13 tumor types. The workflow of this screening process is illustrated in Fig. 4a. Firstly, the DSN-adv model generated a scoring matrix of drug response for patients (PDR) across 13 tumor types, such as COAD with 288 patients (Fi. 4b). Secondly, the top 5% efficient drugs for each patient were filtered out to construct the efficient drugs for patients (EPD) pairs. The histogram plot of EPD pairs revealed that most drugs were inefficient, but a few drugs showed potential efficacy for over half of COAD patients (Fig. 4c and Supplementary Fig. 9). Thirdly, the top 10 drugs with the highest number of efficient patients were selected to form the efficient drugs for indications (EDI) list (Supplementary Table 9). The EDI lists across 13 tumor types were merged to generate the drug efficacy of indication (DEI) table.
Additionally, five C-MDL models were utilized in this workflow to discover drug candidates for comparison. However, the P-SDL models were not applicable in this context as they require patient response data for every drug, which was unavailable.
DEI tables were utilized to verify drug efficacy within specific tumor types. These tables, specific to each of the 13 tumor types, contained information on the efficacy of 17 small molecules, which were considered drug candidates. To gather relevant records for these drug candidates, three databases were accessed: DrugBank[27], ClinicalTrials[28], and Repurposing Hub[29]. Consequently, 10, 9, and 4 drug candidates had corresponding records in these databases, respectively. The information available in DrugBank and ClinicalTrials included tumor type specificity for the drug candidates. This information was then matched with the DEI tables using a heatmap visualization (Fig. 4e,g). Moreover, the ClinicalTrials and Repurposing Hub databases provided additional drug-related information, such as MoA, disease type, clinical phase, research status, etc. Sankey plots were employed to present the relationship between the drug candidates and these research items intuitively and quantitatively (Fig. 4f and Supplementary Fig. 10).
For a fair comparison with previous SOTA models, we defined ”Qualified drugs” as the DEI drugs that met the criteria: i) no false negatives, which means excluding the drugs with any predicted-inefficient EDIs but in-clinical-test recorded EDIs, such as the drug candidate Rapamycin here even though there are some explanations; ii) at least one true positive, which means drugs need to have both predicted-efficient and in-clinical-test EDIs in at least one tumor type. This means that drugs with any predicted inefficient but in-clinical-test recorded EDIs were excluded, ensuring no false negatives. However, false positives were allowed in our analysis. This is because numerous drug-indication combinations have not been tested, and our scenario considers them ”promising”.
Identification of four types of drug sets
DeepCRE were employed to generate the DRP score matrix for the patients of COAD and READ, two subtypes of CRC. This matrix facilitated the ranking of drug candidates based on their efficacy across patient populations, as illustrated in Supplementary Figure 13a. Additionally, we conducted a comprehensive search for clinical records of these drug candidates within the ClinicalTrials database. For clarity and ease of comparison, the evaluation outcomes of DeepCRE alongside traditional methodologies were depicted in an XY plot (Fig. 5c). Notably, the analysis delineated the drug candidates into four distinct sets (A, B, C, and D). Sets A and B comprised drugs favored by DeepCRE predictions but with limited clinical trial evidence, whereas Sets C and D included drugs that were either approved or had substantial clinical trial involvement but were underestimated by DeepCRE predictions.
Experimental assays for the identified drug candidates
To evaluate the efficacy of the identified drug candidates, we employed the In-SMARchip methodology. A total of 16 drug candidates were tested across four concentration gradients (0.2, 2, 20, and 200 uM), with each assay conducted in triplicate. These experiments were performed on eight CRC organoid models. The methodological details, including the experimental setup and analysis, are consistent with those described in our previous work[30, 31, 32].
Docking analysis of small molecules with specific target
The interaction between Sorafenib and the ATP-binding cassette sub-family G member 2 (ABCG2) was meticulously analyzed using Autodock Vina version 1.2.2[48]. This analysis aimed to elucidate the binding poses and molecular interactions critical for the drug’s mechanism of action. These interactions were visualized through the PyMOL software[48], providing a detailed representation of Sorafenib’s binding efficiency and orientation to ABCG2.
Bioinformatics analysis of the pharmacological profile for drug candidates
To discern the pharmacological distinctions between GSK1059615 and other inhibitors targeting the PI3K/mTOR pathway, we extracted perturbation data from the Connectivity Map (CMAP) L1000 database[35]. This dataset was curated to include GSK1059615, WYE-125132 (a comparator PI3K/mTOR inhibitor), and DMSO (as control). Initially, a Principal Component Analysis (PCA) was conducted on the 2,000 most variable genes, followed by a two-dimensional Uniform Manifold Approximation and Projection (Umap) analysis based on the top 50 principal components[49]. Differential expression analysis further identified representative genes, which were visualized on the Umap plot (Fig. 6e). Subsequently, pathway enrichment analysis was performed[50], and the top five up-regulated and down-regulated pathways, determined by adjusted p-values, were illustrated in Fig. 6f, providing insight into the molecular underpinnings and potential therapeutic implications of GSK1059615 compared to other PI3K/mTOR inhibitors.
Code availability
Data availability
DeepCRE dataset, intermediate results, and the final drug response predictions can be found at https://zenodo.org/record/8021167.
Acknowledgements
This research was substantially sponsored by the research projects (Grant No. 2022YFF1203000) supported by the National Key Research and Development Program of China and was substantially supported by the Beijing Academy of Artificial Intelligence (BAAI).
Declaration of Interests
The authors declare no competing interests.